OpenStudy (anonymous):

A bottling machine filled 3 of every 8 bottles with grape juice. The bottling machine filled 144 bottles altogether. How many of these bottles were filled with grape juice?

OpenStudy (greencat):

Total=144

OpenStudy (greencat):

3/8 of Total are grape so: 3/8 of Total=Answer

OpenStudy (greencat):

Let's translate this into math: of means multiply

OpenStudy (greencat):

3/8 of Total=Answer \(\downarrow\) 3/8 * Total=Answer

OpenStudy (greencat):

Can you continue from here?

OpenStudy (anonymous):

yes and thank you @GreenCat

OpenStudy (greencat):

^_^

OpenStudy (anonymous):

can you help me with one more @GreenCat

OpenStudy (greencat):

yes

OpenStudy (anonymous):

The ratio of the perimeters of two rectangles is 5 to 8. The perimeter of the larger rectangle is 72 inches. What is the perimeter of the smaller rectangle? A. 115.2 inches B. 45 inches C. 40 inches D. 16.4 inches

OpenStudy (anonymous):

@GreenCat

OpenStudy (greencat):

Basically, you use the ratio to figure out the perimeter.

OpenStudy (greencat):

The smaller triangle's perimeter should be 5/8 of the larger triangle

OpenStudy (greencat):

Do you understand how I got that?

OpenStudy (anonymous):

but what would that be in decimal form @GreenCat

OpenStudy (greencat):

remebmer that "of" means multiply so 5/8 *72

OpenStudy (greencat):

which would be 45

OpenStudy (anonymous):

thanks @GreenCat

OpenStudy (greencat):

ok bye